OpenStudy (gabylovesyou):

Which equation is a quadratic equation? 5x − 2 = y y + 15x2 = (3x − 2)(5x + 7) y = 3x(5 − x) + 7 y + 2 = (3x2 + 10)(x − 1)

OpenStudy (anonymous):

an equation is quadratic when in standard form y=a x^2 or y=a x^2+b x or y=a x^2+b x +c

Parth (parthkohli):

Or simply, anything in the "degree" (highest power) two with a \(y\) on the left side and \(x\)s on the right side.

Parth (parthkohli):

Yeah, you convert them all such that you get something which looks like \(y = \cdots\).

OpenStudy (anonymous):

isolate for y and they get in a standard form. Solve for y

OpenStudy (anonymous):

example a is already solved for y 5x − 2 = y it isn't a quadratic cuss the x isn't squared

OpenStudy (anonymous):

y = 3x(5 − x) + 7 is also solved for y but not in standard form you need to distribute the 3x to 5 and - x do that and see what happens.

OpenStudy (gabylovesyou):

@HawkCrimson ... help please

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

a quadratic function comes in the form y=a x^2+b x +c where a, b and c r just numbers. so all quadratic functions much have x^2 in it, like look at the first equation 5x − 2 = y. does it have a x^2 in it?

OpenStudy (gabylovesyou):

no it doesnt

OpenStudy (gabylovesyou):

@HawkCrimson

OpenStudy (anonymous):

so it's not a quadratic equation, now lets look at the second one y + 15x2 = (3x − 2)(5x + 7) expand it for me and tell me wat u get?

OpenStudy (gabylovesyou):

y = 11x - 14

OpenStudy (anonymous):

so is it a quadratic equation?

OpenStudy (gabylovesyou):

no

OpenStudy (anonymous):

gud now the 3rd equation y = 3x(5 − x) + 7 expand it, show me it and tell me if it is a quadratic equation

OpenStudy (gabylovesyou):

3(x - 5)x + y = 7

OpenStudy (gabylovesyou):

no it isnt?

OpenStudy (anonymous):

u didnt expand it, u know how to?

OpenStudy (gabylovesyou):

no

OpenStudy (anonymous):

y = 3x(5 − x) + 7 multiply 3x by everything in the brackets which r 5 and -x, do that and show me :)

OpenStudy (gabylovesyou):

15x and 3x^2

OpenStudy (anonymous):

y = 15x − 3x^2 + 7 is it a quadratic formula?

OpenStudy (gabylovesyou):

yes

OpenStudy (gabylovesyou):

thank you! :)

OpenStudy (anonymous):

ur welcome :)